# On the Semantic Latent Space of Diffusion-Based Text-to-Speech Models

### v1

@@ Abstract
Miri Varshavsky Hassid  Roy Hirsch  Regev Cohen  Tomer Golany Daniel FreedmanEhud Rivlin Verily AI {mirivar, royhirsch, regevcohen}@google.com
Preprint under review.
@@ 1 Introduction
Denoising Diffusion Models (DDMs) Sohl-Dickstein et al. (2015) have emerged as a powerful generative tool across a broad variety of tasks and domains. In particular, Text-to-Speech (TTS) systems based on diffusion have shown high-quality speech generation capabilities Huang et al. (2022b); Shen et al. (2023). Although these exhibit improved quality, the extent to which they capture semantic information is yet to be uncovered, and the ability to control the vocal properties (e.g. volume, pitch, gender) of their generated speech is limited. Uncovering the semantic capabilities of TTS diffusion models will allow editing the synthesized speech properties, which is essential in real-world applications, such as human-machine interaction.
@@
Beyond efficiency, researchers have explored DDMs for controllable and expressive TTS. PromptTTS (Guo et al., 2023b) and NaturalSpeech 2 (Shen et al., 2023) employ text prompts and speech prompts, respectively, to control speech style and content. In both methods, the conditional denoiser must undergo a specialized training process. Other methods for controlling the vocal characteristics require large quantities of annotated samples Guo et al. (2023a) or retraining Kim et al. (2022). We propose a speech editing method that requires no additional data or training and can be applied to any frozen diffusion-based TTS model.
In the image synthesis domain, Kwon et al. (2022) recently discovered a semantically meaningful latent space, named h-space, providing versatile semantic editing capabilities. This discovery was further explored by Haas et al. (2023), who proposed methods for identifying semantic directions.
In this work we investigate the existence of a semantic space within diffusion-based TTS systems. We study the properties of h-space in pre-trained TTS models and uncover its acoustically-semantic characteristics. Motivated by this observation, we propose novel methods for semantic speech editing through both supervised and unsupervised latent space arithmetics, inspired by Haas et al. (2023). Our work offers intuitive and efficient audio editing techniques that require neither classifier guidance Guo et al. (2023a), model retraining Kim et al. (2022), optimization, speech prompts nor any architecture modifications. We also present extensive experiments that demonstrate effective and high-quality edited speech synthesis.
@@ 2.1 Denoising Diffusion Models
DDMs generate realistic data by iteratively removing noise, applicable to various modalities like images, audio, and text (Ho et al., 2020). Initially formulated as Markov chains, DDMs can be unified under stochastic differential equations (SDEs) (Song et al., 2020) and adapted for TTS Popov et al. (2021). DDMs consist of two processes: forward diffusion and reverse diffusion. The forward process transforms any data distribution to a Gaussian $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ via an SDE. The reverse diffusion process is defined by another SDE:
@@ 2.2 Semantic Audio Editing via Latent Space Manipulation
Figure 1: The h-space is defined as the concatenation of the bottleneck activations of the U-Net architecture.
@@
Having established the editing framework, we next derive editing directions via the following:
@@ 3.1 Implementation Details
For demonstration, we use Grad-TTS (Popov et al., 2021), a recently proposed publicly available diffusion-based TTS model, trained on LibriTTS (Zen et al., 2019). However, our method can also be applied to any other diffusion TTS model that contains a bottleneck. Grad-TTS takes a text and a speaker embedding as input, and generates a clean mel-spectogram through a U-Net-based denoiser. We use 10 diffusion timesteps for mel-spectogram generation, as suggested by Grad-TTS authors, followed by the Universal HifiGan vocoder (Kong et al., 2020a) for waveform generation.
@@ 3.2 Supervised Latent Space Editing
We begin our analysis by exploring the semantic-capturing capabilities of h-space using the per-speaker gender annotations available for LibriTTS. Following equation 3 we calculate the male-to-female latent direction, and utilize it for audio editing as outlined in equation 2. As the latent vectors’ lengths vary with the input texts, editing direction is defined per text. For a comparable baseline, we use another, simpler, approach for gender-editing: manipulating the speaker embedding, which is provided to the model as an input. We calculate the male-to-female direction in the speaker embedding space by averaging the differences between male and female speakers’ embeddings. The input speaker embedding is modified by adding this direction with different $\lambda$ scales. We provide supplemental samples: https://latent-analysis-grad-tts.github.io/speech-samples/.
Semantic properties evaluation. We fine-tuned a speech gender classifier (Bhamidipati, 2023) on Grad-TTS outputs, acknowledging the different quality of the synthesized speech compared to human-recorded samples. Then, we applied gender editing via both latent space and speaker embedding editing using varying $\lambda$ values, across the first 50 texts of the LibriTTS test set and all 247 speakers. In Figure 2 we report the fraction of samples classified as female, averaged across male and female speakers separately. Latent space editing exhibits a monotonic behavior with more samples classified as female as $\lambda$ increases. On the contrary, speaker embedding editing fails to transform male voices to female ones, and when $\lambda\geq 3$ even originally female voices are not classified as such.
Additionally, 10 human evaluators classified speech samples as male or female. Using samples from 20 different speakers, we compared the unedited Grad-TTS outputs to the gender-edited samples (using $\lambda=2/-2$ for male / female editing, for an effective gender alteration as shown in Figure 2). Table 1 presents the accuracy of predicting the expected gender (original gender for original samples, and contrasting gender for edited samples). Comparing to speaker editing, latent space editing achieves a classification accuracy that is higher by 24%, with statistical significance.
@@
Figure 2: Supervised latent space editing allows gender manipulation, while speaker embedding editing does not. The average percentage of samples classified as female is reported (standard deviation, STD, is shaded).
@@
Table 1: Gender accuracy and MOS (mean $\pm$ STD) for supervised editing. Latent-editing results compared to speaker-editing results are statistically significant (using  Wilcoxon (1945) rank sum test).
Acoustic properties evaluation. To assess the perceived naturalness of the generated speech we measure the Mean Opinion Score (MOS), as quantified by 10 experienced evaluators on a scale of 1 to 5, across the same set of samples reported before. Table 1 shows that the perceived naturalness of latent space editing, compared to speaker editing, is higher by $12\%$, a statistically significant difference. This, combined with the superior perceived gender editing quality, reinforces the latent space’s capability to encapsulate non-trivial semantic information.
@@ 3.3 Unsupervised Latent Space Editing
Next, we investigate semantically meaningful directions in h-space without prior annotations. First, we generated speech samples for the first 50 test texts of LibriTTS and across all 247 speakers, and recorded the latent vectors $h_{T:1}$. Following the unsupervised process defined in Section 2.2, PCA of the latent space was performed for each text across all samples, calculating the first 3 principal components (PCs). As vocal attributes, for each speech sample we extracted its speaker’s gender from the metadata, and measured its intensity, Harmonics-to-Noise Ratio (HNR), and pitch using the Parselmouth Python package (Jadoul et al., 2018).
The latent vectors of each sample were projected onto each PC. Next, we calculated the absolute value Spearman correlation between each vocal attribute and PC-projection vector, averaging across texts and timesteps. As Figure 3 shows, PC1 strongly correlates ($\rho=0.9\pm 0.0$) with speaker’s gender (see Figure 5 in Appendix A), while PC2 correlates ($\rho=0.6\pm 0.1$) with intensity and HNR. Other PCs and vocal attributes show no significant correlation and neither did random projections in the latent space (see Figure 7 in Appendix A).
@@
Semantic properties evaluation. Using PCs as editing directions in h-space, we explore speech editing capabilities. Since the PCs are unitary vectors, the editing directions were normalized to the norm of the latent vectors. Intriguingly, our experiments indicate that decreasing the editing norm at later timesteps improves acoustic quality. As can be seen in Figure 3(a), interpolation along PC1 exhibits a smooth transition between male and female voices. Intensity and HNR decrease when interpolating along PC2 (see Figure 3(b)). Importantly, no gender-editing occurs when interpolating along PC2 (see Figure 6 in Appendix A).
Additionally, we measured the accuracy of gender classification using human annotators and the same 20 speakers. Following the analysis at Figure 4, while editing along PC1 we used $\lambda=3/-3$ for male/female speakers, for effective gender alteration. For PC2, $\lambda=-2$ was used to maximize HNR. PC1-edited samples were successfully classified as the contrasting gender with an even higher accuracy than un-edited ones (Table 2).
@@
Figure 4: Interpolating along the semantic directions revealed by PCA changes the vocal attributes accordingly. Shaded area is the standard deviation.
@@
Acoustic properties evaluation. Using the same setup, we assessed speech naturalness using MOS. Table 2 compares the perceived naturalness of samples with and without latent editing. The Wilcoxon rank sum test indicated no statistically significant difference in the MOS between groups (p-value $\gg 0.05$). Thus, we conclude that speech editing through unsupervised latent space manipulation does not compromise the acoustic quality.
@@ 4 Conclusions
In this paper, we identify the semantic properties of the latent space of diffusion-based TTS models, referred to as h-space. We develop supervised and unsupervised methods for finding interpretable directions in that space, and provide empirical qualitative evidence for their semantic quality. Moreover, the proposed latent space editing methods preserve and even enhance the acoustic quality of the generated samples. This study presents evidence regarding specific vocal attribute manipulation, such as gender or intensity. However, the presented method can be applied to any vocal attribute.
@@
This study is subject to several limitations. We demonstrated our analysis on the Grad-TTS model Popov et al. (2021) (trained on LibriTTS dataset (Zen et al., 2019)), and used the Universal HifiGAN (Kong et al., 2020a) for waveform generation. We do not develop novel TTS models from scratch, and focus on analysing existing systems. These are all publicly available for our research purposes. Under these settings, several limitations apply to our analysis:
@@ B.2 Mean Opinion Score Evaluation
Score Qaulity
5.0 Excellent (Completely undefined)
@@
4.0 Good (Mostly undefined
@@
3.0 Fair (Fair (Equally undefined and undefined))

### v2

@@ Abstract
Miri Varshavsky-Hassid  Roy Hirsch  Regev Cohen  Tomer Golany Daniel FreedmanEhud Rivlin Verily AI {mirivar, royhirsch, regevcohen}@google.com
**footnotetext: Equal contributionPaper accepted to ACL 2024.
@@ 1 Introduction
Denoising Diffusion Models (DDMs) Sohl-Dickstein et al. (2015) have emerged as a powerful generative tool across a broad variety of tasks and domains. In particular, Text-to-Speech (TTS) systems based on diffusion have shown high-quality speech generation capabilities Huang et al. (2022b); Shen et al. (2023). Although these exhibit improved quality, the extent to which they capture semantic information is yet to be uncovered, and the ability to control the vocal properties (e.g. volume, pitch, gender) of their generated speech is limited. Uncovering the semantic capabilities of TTS diffusion models will allow editing the properties of synthesized speech, which is essential in real-world applications, such as human-machine interaction.
@@
Beyond efficiency, researchers have explored DDMs for controllable and expressive TTS. PromptTTS (Guo et al., 2023b) and NaturalSpeech 2 (Shen et al., 2023) employ text prompts and speech prompts, respectively, to control speech style and content. In both methods, the conditional denoiser must undergo a specialized training process. Other methods for controlling the vocal characteristics require large quantities of annotated samples Guo et al. (2023a) or retraining Kim et al. (2022). We propose a speech editing method that requires no additional data or training and can be applied to any frozen diffusion-based TTS model that incorporates a bottleneck.
In the image synthesis domain, Kwon et al. (2022) recently discovered a semantically meaningful latent space, named h-space, providing versatile semantic editing capabilities. This discovery was further explored by Haas et al. (2023), who proposed methods for identifying semantic directions. To the best of our knowledge, despite the widespread adoption of diffusion models for TTS in recent years, the existence of a hidden semantic space has not been examined in the speech synthesis domain. This raises intriguing questions regarding the possibility of facilitating latent space arithmetics for audio editing.
In this work we investigate the existence of a semantic space within diffusion-based TTS systems. We study the properties of h-space in pre-trained TTS models and uncover its acoustically-semantic characteristics. Then, we propose novel methods for semantic speech editing through both supervised and unsupervised latent space arithmetics, inspired by Haas et al. (2023) and adapted to the speech synthesis domain for the first time. Our work offers intuitive and efficient audio editing techniques that require neither classifier guidance Guo et al. (2023a), model retraining Kim et al. (2022), optimization, speech prompts nor any architecture modifications. To validate our methods, we present extensive experiments that demonstrate effective and high-quality edited speech synthesis.
@@ 2.1 Denoising Diffusion Models
DDMs generate realistic data by iteratively removing noise, and are applicable to various modalities like images, audio, and text (Ho et al., 2020). Initially formulated as Markov chains, DDMs can be unified under stochastic differential equations (SDEs) (Song et al., 2020) and adapted for TTS Popov et al. (2021). DDMs consist of two processes: forward diffusion and reverse diffusion. The forward process transforms any data distribution to a Gaussian $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ via an SDE. The reverse diffusion process is defined by another SDE:
@@ 2.2 Semantic Audio Editing via Latent Space Manipulation
Figure 1: The h-space of a diffusion model is defined as the concatenation of the bottleneck activations of the U-Net architecture.
@@
Figure 2: We propose a simple yet effective semantic audio-editing method. A latent semantic direction is defined either in a supervised or an unsupervised manner, and the corresponding speech attribute is edited by applying that direction to the latent space during the generation process of a new speech sample. The method is demonstrated with the male-to-female editing direction.
Having established the editing framework, we next derive editing directions via the following (illustrated in Figure 2):
@@ 3.1 Implementation Details
For demonstration, we use Grad-TTS (Popov et al., 2021), a recently published publicly available diffusion-based TTS model, trained on LibriTTS (Zen et al., 2019). However, our method can also be applied to any other unguided diffusion-based TTS model that contains a bottleneck. Grad-TTS takes a text and a speaker embedding as input, and generates a clean mel-spectogram through a U-Net-based denoiser. We use 10 diffusion timesteps for mel-spectogram generation, as suggested by Grad-TTS authors, followed by the Universal HifiGan vocoder (Kong et al., 2020a) for waveform generation.
@@ 3.2 Supervised Latent Space Editing
We begin our analysis by exploring the semantic-capturing capabilities of h-space using the per-speaker gender annotations available for LibriTTS. Capturing the latent code during all timesteps of the generation process and following Equation 3, we calculate the male-to-female latent direction, and utilize it for audio editing as outlined in Equation 2. As the latent vectors’ lengths vary with the input texts, editing direction is defined per text. For a comparable baseline, we use another, simpler, approach for gender-editing: manipulating the speaker embedding, which is provided to the model as an input. We calculate the male-to-female direction in the speaker embedding space in a similar manner by averaging the differences of speaker embeddings between pairs of male and female speakers. The input speaker embedding is modified by adding this direction with different scales ($\lambda$). We provide supplemental samples, demonstrating the suggested audio editing methods: https://latent-analysis-grad-tts.github.io/speech-samples/.
Semantic properties evaluation. We fine-tuned a speech gender classifier (Bhamidipati, 2023) on Grad-TTS outputs, acknowledging the different quality of synthesized speech compared to human-recorded samples. Then, we applied gender editing via both latent space and speaker embedding editing using varying $\lambda$ values, across the first 50 texts of the LibriTTS test set and all 247 speakers. In Figure 3 we report the fraction of samples classified as female for each $\lambda$ value, averaged across input male and female speakers separately. Latent space editing exhibits a monotonic behavior with more samples classified as female as $\lambda$ increases. On the contrary, speaker embedding editing fails to transform male voices to female ones, and when $\lambda\geq 3$ even originally female voices are not classified as such.
Additionally, 10 human evaluators classified speech samples as male or female. Analyzing samples from 20 different speakers, we compared the unedited Grad-TTS outputs to the gender-edited samples. For an effective gender alteration as shown in Figure 3, we used $\lambda=2$ and $-2$ for male-to-female and female-to-male editing, respectively. Table 1 presents the accuracy of predicting the expected gender (original gender for original samples, and contrasting gender for edited samples). Comparing to speaker editing, latent space editing achieves a classification accuracy that is higher by 24%, with statistical significance (p-value $<0.001$).
@@
Figure 3: Supervised latent space editing allows gender manipulation, while speaker embedding editing does not. The percentage of samples classified as female is reported separately for male and female input spakers, averaged across 50 texts and all speakers (standard deviation, STD, is shaded).
@@
Table 1: Supervised latent space editing generates intelligible samples where the perceived speaker’s gender is correctly classified, while speaker embedding editing does not. Average gender accuracy and MOS (mean $\pm$ STD) are reported. Latent-editing results compared to speaker-editing results are statistically significant (using  Wilcoxon (1945) rank sum test).
Acoustic properties evaluation. To assess the perceived naturalness of the generated speech we measure the Mean Opinion Score (MOS), as quantified by 10 experienced evaluators on a scale of 1 to 5, across the same set of samples reported before. Table 1 shows that the perceived naturalness of latent space editing, compared to speaker editing, is higher by $12\%$, a statistically significant difference (p-value $<0.001$). This, combined with the superior perceived gender editing quality, reinforces the latent space’s capability to encapsulate non-trivial semantic information.
@@ 3.3 Unsupervised Latent Space Editing
Next, we investigate semantically meaningful directions in h-space without prior annotations. First, we generated speech samples for the first 50 test texts of LibriTTS and across all 247 speakers, and recorded the latent vectors $\textbf{h}_{T:1}$. Then, following the unsupervised process defined in Section 2.2, PCA of the latent space was performed for each text across all samples, calculating the first 3 principal components (PCs). As vocal attributes, for each speech sample we extracted its speaker’s gender from the metadata, and measured its intensity, Harmonics-to-Noise Ratio (HNR), and pitch using the Parselmouth Python package (Jadoul et al., 2018).
The latent vectors of each sample were projected onto each PC. Next, we calculated the absolute value Spearman correlation between each vocal attribute and PC-projection vector, averaging across texts and timesteps. As Figure 4 shows, PC1 strongly correlates ($\rho=0.9\pm 0.0$) with speaker’s gender (also see Figure 6 in Appendix A), while PC2 correlates ($\rho=0.6\pm 0.1$) with intensity and HNR. Other PCs and vocal attributes show no significant correlation and neither did random projections in the latent space (see Figure 8 in Appendix A).
@@
Semantic properties evaluation. Using PCs as editing directions in h-space, we explore speech editing capabilities. Since the PCs are unitary vectors, the editing directions were normalized to the norm of the latent vectors. Intriguingly, our experiments indicate that decreasing the editing norm at later timesteps improves acoustic quality. As can be seen in Figure 5(a), interpolation along PC1 exhibits a smooth transition between male and female voices. Simialrly, intensity and HNR decrease when interpolating along PC2 (see Figure 5(b)). Importantly, no gender-editing occurs when interpolating along PC2 (see Figure 7 in Appendix A).
Additionally, we measured the accuracy of gender classification as evaluated by human annotators on the same 20 speakers. Following the analysis in Figure 5, to ensure effective gender alteration, we used $\lambda=3$ or $-3$ for originally male or originally female speakers, respectively, while editing along PC1. For PC2, $\lambda=-2$ was used to maximize HNR. PC1-edited samples were successfully classified as the contrasting gender with an even higher accuracy than un-edited ones (Table 2).
@@
Figure 5: Interpolation along the semantic directions revealed by PCA changes the vocal attributes accordingly. The reported values are averaged over 50 texts and all speakers. Shaded area is the STD.
@@
Acoustic properties evaluation. Using the same setup, we assessed speech naturalness using MOS. Table 2 compares the perceived naturalness of samples with and without latent editing, presenting similar scores between the groups. The Wilcoxon rank sum test indicated no statistically significant difference in the MOS between groups (p-value $\gg 0.05$). Thus, we conclude that speech editing through unsupervised latent space manipulation does not compromise the acoustic quality.
@@ 4 Conclusions
In this paper, we identify the semantic properties of the latent space of diffusion-based TTS models, referred to as h-space. We develop supervised and unsupervised methods for finding interpretable directions in that space, and provide empirical qualitative evidence for their semantic quality. Moreover, the proposed latent space editing methods preserve and even enhance the acoustic quality of the generated samples. This study presents evidence regarding specific vocal attribute manipulation, such as gender or intensity. However, the presented method can be applied to any vocal attribute present in the data.
@@
This study is subject to several limitations. We demonstrated our analysis on the Grad-TTS model Popov et al. (2021) (trained on LibriTTS dataset (Zen et al., 2019)), and used the Universal HifiGAN (Kong et al., 2020a) for waveform generation. These are all publicly available for our research purposes. We do not develop novel TTS models from scratch, and focus on analysing existing ones. Under these settings, several limitations apply to our analysis:
@@
    Our method is general and can be applied to any frozen unguided diffusion-based TTS model that contains a bottleneck. However, since we were limited to publicly available models, we chose to focus on analysing the Grad-TTS model.
  4. 4.
@@ B.2 Mean Opinion Score Evaluation
Score Quality
5.0 Excellent (Completely defined)
@@
4.0 Good (Mostly defined)
@@
3.0 Fair (Equally defined and undefined)